Need help with the rest and how it was gotten
Parts E and F I dont understand how to solve for them. the front page has all the data collected and what i was able to solve for
E. Molar Volume of CO; Gas Trial I {nL 1.030 1. Pressure of dry CO,(g) (atm). See Part C.7 2. Volume of CO-(g) at STP (L). See equation 13.4. ulayanloM ebixo 3. Molar volume of CO,(g) at STP (Lmol). See equation 13.2. 4. Average molar volume of CO (g) at STP (L/mol) ae ab sa27.4jit F.Percent CaCO, in Mixture 1. Moles of CaCO, in sample from mol CO, generated (mol) See equation 13.1. 2. Mass of CaCO, in sample (g) 3. Mass of original sample (g) 4. Percent of CaCo, in sample (%) 5. Average percent of CaCO, in sample (%) Laboratory Questions Circle the questions that have been assigned 1. Part A.1. The water for the pan (Part A.3) is not saturated with CO2 Will the reported percer high, too low, or unaffected? Explain. sample be too 2. Part A.2. Suppose 6 M HCI is substituted for the 3 M HCl in the procedure. What would be substitution? Explain.
Experiment 13 Report Shee A Carbonate Analysis Molar Volume of Carbon Dioxide Desk No. Name Lab Sec Date A. Sample Preparation and Setup Apparatus Calculation of mass of CaCO, sample for analysis (from Prelaboratory Assignment, question 2). Ca (itcoz )2 + 2 HCI -Hz02 Coz tCacir t 2 SI A Trial 2 Trial I Unknown sample no. O.191ym 63.443m 1. Mass of sample (g) Mass of generator+ sample before reaction (e) 2. 3. Instructor's approval of apparatus C. Determination of Volume, Temperature, and Pressure of the Carbon Dioxide Gas 1. Initial reading of volume of water in CO,-collecting graduated cylinder (mL) 2. Final reading of volume of water in CO-collecting graduated cylinder (mL) 11m L 30ML Volume of CO-(R) collected (L) 3. oht 7748.53870 4. Temperature of water (C) Conet to 5. Barometric pressure (torr) 22.4od fremegine c (torr) Vapor pressure of H,O at 6 6.13810 7. Pressure of dry CO,(g) (torr) 141.535 22- 7 D. Amount of Carbon Dioxide Gas Evolved 63.342 1. Mass of generator + sample after reaction (g) 2. Mass loss of generator mass CO, evolved (g) 3. Moles of CO, evolved (mol) mel 1:n Hy 25-4 7orr 1q.47in Ms 748-5 3g Tod Experiment 13 9.101uu.01 0. 0 072
Answer #1

Pressure of dry CO2 , P1 = 726.138 torr = 0.955 atm ( 760 torr = 1 atm)

Volume of CO2 collected , V1= 30 ml = 30/1000 = 0.03 lts

Temperature , T1 = 24 +273 = 297 K

P2 = 1 atm (STP)

T2 = 273 K (STP)

Volume of CO2 at STP , V2 (P1V1/T1 = P2V2/T2) = 0.955 atm x 0.03 lts x 273 K/297 K X 1 atm = 0.026 lts

Moles of CO2 generated , n = PV/RT = 0.955 atm x 0.026 lts/0.0821 lt atm/mol K x 297 K = 0.02483/24.38 = 0.001 moles

Molar volume of CO2 at STP = 0.026 lt/0.001 moles = 26 l/mol

CaCO3 + 2HCl ---> CaCl2 + CO2(g)

Molar ratio of CaCO3 and CO2 = 1:1

So, moles of CaCO3 = moles of CO2 = 0.001 moles

Molar mass of CaCO3 = 100 g/mol

Mass of CaCO3 = moles x molar mass = 0.001 mol x 100 g/mol = 0.1 g

Mass of original sample = 0.191 g

Percent of CaCO3 in the original mixture = (practical mass/original mass) x 100 = (0.1 g/0.191 g) x 100 = 52.36 %